OpenStudy (meilendurcer):

Select the equations that are parallel and perpendicular to y = x + 5 and that pass through the point (-2, -1). A. parallel: y = -x - 1 perpendicular: y = x + 2 B. parallel: y = x - 1 perpendicular: y = -x + 1 C. parallel: y = x + 1 perpendicular: y = -x - 3 D. parallel: y = 2x - 2 perpendicular: y = -2x - 1

OpenStudy (anonymous):

The coefficient of 'x' in y = mx + b form of an equation gives you the slope of the equation. So the slope of the equation y = x + 5 is 1. Parallel lines have the same slope, so in each of the choices given we can expect 'x' to have an coefficient of 1, that is, an equation of the form y = x + something Perpendicular lines have slopes whose product is -1. So if 1 is the slope of a line, then the slope of the line perpendicular to it must be -1 since 1 * -1 = -1. So in each of the choices given you would expect 'x' to have a coefficient of -1, that is, an equation of the form y = -x + something

OpenStudy (meilendurcer):

@navk could the answer be D?

OpenStudy (meilendurcer):

@navk please check my answer

OpenStudy (anonymous):

It's not D. From the parallel and perpendicular part we know that it can only be B or C since a parallel equation to y = x + 5 will be like 'y = x + something' while a perpendicular to it will be like y = -x + something

OpenStudy (anonymous):

It is also given that the equations pass through the point (-2, -1) so try plugging in x = -2 and y = -1 into the options in B and C to get the correct choice

OpenStudy (meilendurcer):

how do i do that? i can't really do this type of math